MelSaverethiar:

Prove that the sequence {an} defined by the relation a1 = 1, an = 1 +1/1! +1/2! + ... +1/(n − 1)! converges.

YRJ8498:

Is this a Quiz or a test?

MelSaverethiar:

its a question on my assigment

YRJ8498:

So what do you know so fat about this lesson?

MelSaverethiar:

what just happened?

MelSaverethiar:

bounded below...?

YRJ8498:

So what do you know so far about this lesson Mel?

MelSaverethiar:

idk

YRJ8498:

you have to work with me so i can help so please tell me what u understand and don't

MelSaverethiar:

i dont undersand the whole chapter

yy:

use the questions one

MelSaverethiar:

what does it mean to be bounded above or below ?

MelSaverethiar:

ok thank u

surjithayer:

\[\frac{ a_{n} }{ a_{n-1} }=\frac{ \frac{ 1 }{ (n-1)! } }{\frac{ 1 }{ (n-2)! } }=\frac{ (n-2)! }{ (n-1)! }\] \[=\frac{ (n-2) }{ (n-1)(n-2)! }=\frac{ 1 }{ n-1 }<1\] so it is a decreasing function. also \[\lim_{n \rightarrow \infty}\frac{ 1 }{ n-1 }\rightarrow 0\] so it is convergent.